40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Feb 13 13:35:38 2009, Yoichi, Update, LSC, Locking status 
    Reply  Sat Feb 14 16:53:26 2009, rob, Update, LSC, Locking status 
       Reply  Sun Feb 15 09:35:00 2009, Yoichi, Update, LSC, Locking status 
       Reply  Sun Feb 15 15:53:21 2009, Rob, Update, LSC, Locking status 
          Reply  Mon Feb 16 10:18:13 2009, Alberto, Update, LSC, FE system rebooted 
             Reply  Mon Feb 16 14:12:21 2009, Yoichi, Update, LSC, FE system rebooted 
Message ID: 1306     Entry time: Sun Feb 15 15:53:21 2009     In reply to: 1304     Reply to this: 1308
Author: Rob 
Type: Update 
Category: LSC 
Subject: Locking status 

Quote:

I didn't get a chance to do much testing since the sus controller (susvme1) went nuts. In retrospect, this could be due to something in the script, so maybe we should try a burt restore to Friday afternoon next time someone wants to look at it.


I tried the burt restore today, it didn't work. Also tried some switching of timing cables, and multiple reboots, to no avail. This will require some more debugging. We might try diagnosing the clock driver and fanout modules, the penteks, and we can also try rebooting the whole FE system.
ELOG V3.1.3-